OpenStudy (jjuden):

Divide. (15mn-20n)/(5n)

OpenStudy (anonymous):

Common factor : n

OpenStudy (jjuden):

\[\frac{ 15mn-20n }{ 5n }\]

OpenStudy (anonymous):

Again, this can be written as: \[\frac{ 15mn}{ 5n } - \frac{20n}{5n}\] Which will give you: \[3m - 4\]

OpenStudy (mathmale):

Actually, the common factor (greatest common factor) of the numerator is 5n. @jjuden: Please factor 5n out of the numerator. Then simplifiy the resulting ratio.